From GlueXWiki
Jump to: navigation, search

March 15, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
    • Testing at Blue Crab (Lubomir)
  2. Oxygen contamination in the first package FDC ELOG (Beni, Bill, Lubomir)
  3. Engineering update (Bill)
  4. Electronics update (Chris)
  5. Other


Participants: Dave, Bill, John, Nick, Eugene, Beni, Simon, and Lubomir.


- Dave for the third package: wire planes #4 and #13 remain to be deadened. The chemical re-crystalized and has to be reheated. Chris will put the components on #13 on Friday. Mylar will be tensioned on cathode #40 type-2 and we will be done with the parts for the third package. Package four: started working on the first two cathodes type-3, laminating and testing the wire frames. Most likely Anatoly will start working on BCAL lightguides gluing (in couple of weeks?), so we would like him first to replace the HV caps and do the testing on the rest of the wire frames.

- Lubomir: third cell on the third package had to be opened twice to clean a noise place (10-20 kHz high amplitude) on a sense wire; turned out to be a small hair on the back side of the wire. Fourth cell was installed too but not tested yet. As Eugene suggested on the tracking meeting we discussed if we stop the package assembly till we figure out the oxygen issue. Everybody agreed. We will continue mostly with the cathode production and wire frame production, preferably without stringing the wires to avoid damage/cleaning.

Oxygen contamination in the first package

- Beni showed different results from the FDC ELog about efficiency dependence on the gas flow for three different Ar/CO2 gas mixtures we had recently. The efficiency is between 90-100% with 90/10 at high flow (400-500 scc/min). For 40/60 at 500 scc/min the efficiency is 60-70% but very uniform. Beni studied also the dependence of the oxygen percentage on the flow and differential pressure in the chamber. Actually the two parameters were varied at the same time, but by using a pump to lower the pressure almost to zero Pa, Beni didn't see strong dependence on the pressure. For 500 scc/min and ~150Pa the best value for the oxygen was 6200vpm, while with 300scc/min and ~250Pa oxygen went up to ~8000vpm.

- Casey started replacing the plastic rods with aluminum ones on the first package and applying 100 in*oz torque instead of 50; not an easy work since some of the plastic broke. Bill explained that with 50 in*oz two surfaces must be flush when the O-ring has its max compression of 35%. Due to possible deformation this may not be the case and increasing the torque will just make the frames flatter. Higher torque will be applied on the test chamber as well and look at the oxygen percentage.

- Future plans: we will start flushing the second package with a premix 40/60 bottle at Blue Crab and later measure the oxygen there. We discussed again if we can use other gases like He or Ne to test for leakage. If changing the rods doesn't fix the problem we will consider replacing the O-rings. Bill ordered new O-rings: Buna-N (Nitrile) and according to a table where this material is mentioned with a similar name, the oxygen permeability must be an order of magnitude better. First we will change the O-rings in the testing chamber.


- Bill working now on the package connections and cooling system drawings.

- We discussed if the Fluorinert will be a problem for our leaky chambers. It is known that very small percentge of Fluorine can cause serious problems in the chamber, but what about Fluorinert? Eugene suggested talking to experts.


- Nick: 5 additional pre-amp sets (22 each) ready, plus one at Blue Crab, we have the sets needed for the second package.

- Nick made the control for the fan pump that is installed now in an Al cylinder that Bill designed for it, ready to be tested on the first package.